How many bit strings of length 8 either start with 1 bit or end with 2 bits 0
1
Expert's answer
2020-11-02T20:45:54-0500

"\text {let }A=\{1xxxxxxx\}\text{where }x\isin\{0,1\}"

"\vert{A}\vert=2^7=128"

"\text {let }B=\{xxxxxx00\}\text{where }x\isin\{0,1\}"

"\vert{B}\vert=2^6=64"

"A\cap{}B=\{1xxxxx00\}\text{where }x\isin\{0,1\}"

"\vert{A\cap{}B}\vert=2^5=32"

"\text{according to Inclusion-exclusion principle }\vert{A\cup{}B}\vert=\vert{A}\vert+\vert{B}\vert-\vert{A\cap{}B}\vert"

"\vert{A\cup{}B}\vert=128+64-32=160"

Answer: 160